# Deterministic Dynamic Maximal Matching in Sublinear Update Time

### v1

@@ 56             Delete-$M_{\texttt{adj}}$($(u,v)$)
Suppose that we maintain the matching $M_{\texttt{base}}$ as per Lemma 6.6. Then with an additive overhead of $\tilde{O}\left(\delta n+\frac{B}{\delta}+\frac{n}{\delta B}\right)$ amortized update time, we can deterministically and explicitly maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$.
We devote the rest of Section 6.4 to the proof of Lemma 6.14. We start by summarizing a few key properties of the adjunct graph $G_{\texttt{adj}}$ in the claim below.
@@
Consider any node $v\in V_{\texttt{adj}}\setminus V_{\texttt{dmg}}$. By part-(1) of Claim 6.15, we have $v\notin V_{\texttt{sf}}$, and hence $v\notin V_{\texttt{hi}}$ (see Definition 6.2). As the node-set $V$ is partitioned into the subsets $V_{\texttt{lo}},V_{\texttt{med}},V_{\texttt{hi}}$ (see Section 6.2), it follows that $v\in V_{\texttt{lo}}\cup V_{\texttt{med}}$.
Since each phase lasts for $\delta n$ updates in $G$, we infer that $|E\setminus E_{\texttt{init}}|\leq\delta n$ (see Section 6.2). Thus, the node $v$ can have at most $\delta n$ incident edges in $G_{\texttt{adj}}$ that are part of $E\setminus E_{\texttt{init}}$. Henceforth, we focus on bounding the number of incident edges on $v$ in $G_{res}$ that are part of $E\cap E_{\texttt{init}}$. We start by partitioning this concerned set of edges into three subsets, depending on their other endpoints:
$Z_{\texttt{lo}}$ ::=:= $\{(u,v)\in E_{\texttt{adj}}\cap E_{\texttt{init}}:u\in V_{\texttt{lo}}\cap V_{\texttt{adj}}\},$
$Z_{\texttt{med}}$ ::=:= $\{(u,v)\in E_{\texttt{adj}}\cap E_{\texttt{init}}:u\in V_{\texttt{med}}\cap V_{\texttt{adj}}\},$
$Z_{\texttt{hi}}$ ::=:= $\{(u,v)\in E_{\texttt{adj}}\cap E_{\texttt{init}}:u\in V_{\texttt{hi}}\cap V_{\texttt{adj}}\}.$
By part-(3) of Proposition 6.1, we have $Z_{\texttt{lo}}\subseteq E(H_{\texttt{init}})$. Since $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it has maximum degree $B$ (see Definition 2.1), and hence we get $|Z_{\texttt{lo}}|\leq B$. Next, part-(3) of Claim 6.15 implies that $|Z_{\texttt{med}}|\leq 22\delta n=O(\delta n)$. Finally, since $V=V_{\texttt{sf}}\cup V_{\texttt{dmg}}$ (see Definition 6.2), by part-(1) and part-(2), we get $|Z_{\texttt{hi}}|=O\left(\frac{n}{B}\right)$.
@@
At the start of a phase, we have $V_{\texttt{adj}}\cap V_{\texttt{hi}}=(V_{\texttt{adj}}\cap V_{\texttt{sf}})\cup(V_{\texttt{adj}}\cap V_{\texttt{dmg}})=\emptyset$, by Definition 6.2 and part-(1) and part-(2) of Claim 6.15. Accordingly, at this point in time, the node-set $V_{\texttt{adj}}$ is partitioned into two subsets: $V_{\texttt{adj}}\cap V_{\texttt{lo}}$ and $V_{\texttt{adj}}\cap V_{\texttt{med}}$.
By part-(3) of Claim 6.15, there are at most $O(\delta n\cdot\delta n)$ edges in $E_{\texttt{adj}}$ whose both endpoints are in $V_{\texttt{adj}}\cap V_{\texttt{med}}$.
Next, consider any node $u\in V_{\texttt{adj}}\cap V_{\texttt{lo}}$. Since $V_{\texttt{adj}}\cap V_{\texttt{hi}}=\emptyset$ at this point in time, every edge in $E_{\texttt{adj}}$ that is incident on $u$ has its other endpoint in $V_{\texttt{lo}}\cup V_{\texttt{med}}$ (see Section 6.2). Accordingly, by part-(3) of Proposition 6.1, every edge in $E_{\texttt{adj}}\subseteq E_{\texttt{init}}$ that is incident on $u$ is part of $H_{\texttt{init}}$. Since $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it has maximum degree $B$ (see Definition 2.1). Thus, the node $u\in V_{\texttt{adj}}\cup V_{\texttt{hi}}$ has degree at most $B$ in $G_{\texttt{adj}}$, at this point in time. We conclude that there are at most $O(nB)$ edges in $E_{\texttt{adj}}$ that have some endpoint in $V_{\texttt{adj}}\cap V_{\texttt{lo}}$.
@@
6.4.1 Proof of Lemma 6.14.
At the start of a phase, we initialize $M_{\texttt{adj}}$ to be any arbitrary maximal matching in $G_{\texttt{adj}}$. By Corollary 6.17, this takes $O(\delta n\cdot\delta n+Bn)$ time. Since the phase lasts for $\delta n$ updates in $G$, this step incurs an amortized update time of:
@@
Subsequently, to handle the updates during a phase, we use the following auxiliary data structures: Each node $v\in V_{\texttt{dmg}}$ maintains the set $\mathcal{F}_{\texttt{adj}}(v):=\left\{u\in V_{\texttt{adj}}\setminus V(M_{\texttt{adj}}):(u,v)\in E_{\texttt{adj}}\right\}$ of its free neighbors (in $G_{\texttt{adj}}$) under the matching $M_{\texttt{adj}}$. [^8] Maximality of $M_{\texttt{adj}}$ implies that $\mathcal{F}_{\texttt{adj}}(v)=\emptyset$ for all $v\in V_{\texttt{dmg}}\setminus V(M_{\texttt{adj}})$. Whenever a node $v$ moves from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$, we spend $\tilde{O}(n)$ time to initialize the set $\mathcal{F}_{\texttt{adj}}(v)$ as a balanced search tree. Within a phase, a node can move from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$ at most once (see Definition 6.2 and Corollary 6.3). Thus, by part-(2) of Claim 6.15, these initializations take $\tilde{O}\left(\frac{n}{B}\cdot n\right)=\tilde{O}\left(\frac{n^{2}}{B}\right)$ total time during a phase. As each phase lasts for $\delta n$ updates in $G$, this incurs an amortized update time of $\tilde{O}\left(\frac{n^{2}}{B\cdot\delta n}\right)=\tilde{O}\left(\frac{n}{\delta B}\right)$.
Since $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$, Condition 6.5 and Lemma 6.6 imply that each edge update in $G$ leads to $O(1)$ node-updates in $G_{\texttt{adj}}$. Moreover, while maintaining the maximal matching $M_{\texttt{adj}}$ in $G_{\texttt{adj}}$, a node-update can be handled in $\tilde{O}\left(\delta n+B+\frac{n}{B}\right)$ worst-case time, for the following reason.
Suppose that we are trying to find a free neighbor (if it exists) of some node $v\in V_{\texttt{adj}}$. This can be done in $\mathrm{deg}_{G_{\texttt{adj}}}(v)=O\left(\delta n+B+\frac{n}{B}\right)$ time if $v\in V_{\texttt{adj}}\setminus V_{\texttt{dmg}}$ (see Corollary 6.16), and in $\tilde{O}(1)$ time if $v\in V_{\texttt{dmg}}$ (using the set $\mathcal{F}_{\texttt{adj}}(v)$). Further, once we decide to match a node $u\in V_{\texttt{adj}}$ to one of its free neighbors, we can accordingly update all the relevant sets $\mathcal{F}_{\texttt{adj}}(v)$ in $\tilde{O}\left(\left|V_{\texttt{dmg}}\right|\right)=\tilde{O}\left(\frac{n}{B}\right)$ time (see part-(2) of Claim 6.15).
@@
Lemma 6.14 follows from (13) and (14).
@@
Theorem 1.1 follows from Lemmas 6.6 and 6.14.
To be a bit more specific, first recall that maintaining the EDCS $H$ incurs an update time of $O\left(\frac{n}{\epsilon B}\right)=O\left(\frac{n}{\delta B}\right)$, according to (6) and the first paragraph of Section 6.1. Now, for the deterministic algorithm, part-(1) of Lemma 6.6 incurs an amortized update time of $\tilde{O}\left(B\cdot n^{1/2}\cdot\delta^{-3/2}\right)$. Thus, by Lemma 6.14, the overall amortized update time becomes
@@
Finally, for the randomized algorithm, part-(2) of Lemma 6.6 incurs an amortized update time of $\tilde{O}(B\delta^{-1}+\delta^{-3})$. Thus, by Lemma 6.14, the overall amortized update time becomes
@@ 7 Open Problems
The algorithm from Lemma 5.1 is the main bottleneck in our dynamic maximal matching algorithm. In the static setting, we can deterministically find a perfect matching in an $m$-edge graph with $\gamma$-degree-gap with $\tilde{O}(m/\gamma)$ time, for example, using Dinic’s flow algorithm. Can we deterministically match this running time in the decremental setting? Such an algorithm would immediately imply a deterministic dynamic maximal matching algorithm with $\tilde{O}(n^{4/5})$ update time via our framework.
@@ Appendix A Convergence of Random Walks in Eulerian Expanders
In this section, we prove Theorem 5.6 for completeness. We follow the notations from [RTV06]. We say that a matrix $M\in[0,1]^{V\times V}$ is a Markov chain with vertex set $V$ if $\sum_{v\in V}M(v,u)=1$ for every vertex $u$. Given a directed graph $G$, the random walk matrix $M_{G}$ of $G$ is the matrix where, for each entry $(v,u)$,
@@
For an Eulerian graph $G$, observe that the conductance of $M_{G}$ defined above coincides with the conductance of $G$ defined in Equation 4. That is, we have
@@
With all lemmas above, we are now ready to prove Theorem 5.6.
Proof of Theorem 5.6.
@@
Let $\alpha$ be the distribution over $V$ where $\alpha(u)=1$ if $u=v$ and $\alpha(u)=0$ for $u\neq v$. Observe that $M_{G}^{k}\alpha$ is precisely the distribution over $V$ after making $k$ random walk steps in $G$ starting from vertex $v$. So $p$ is precisely the $t$-th entry of $M_{G}^{k}\alpha$. By Lemma A.3, we have that
@@
To bound $\lambda_{\pi}(M_{G})$, observe that $M_{G}(u,u)\geq 1/2$ for every $u$ because the number of self loops at $u$ is at least $\mathrm{deg}(u)/2$. So, by Lemma A.4 and Equation 15, we have
@@
Appendix B Proof of Lemma 3.1
@@
Now, we are ready to prove Lemma 3.1.
@@
Note that when the algorithm of Theorem B.1 is called on $G^{\prime}$ in $O(|V^{\prime}|\log|V^{\prime}|)=O(n\log n)$ expected time it returns a matching $M$ such that every vertex in $V_{\kappa}$ is not matched by $M^{\prime}$ with probability at most $2/\Delta$. Therefore, the expected number of unmatched vertices in $V_{\kappa}$ under $M$ is at most $22|V_{\kappa}|/\Delta\leq 2n/\Delta\leq n\kappa/$ 2. Define a matching $M$ returned by the algorithm to be suitable if it doesn’t match at most $n\kappa$ vertices of $V_{\kappa}$. By Markov’s inequality we can argue that the algorithm of Theorem B.1 returns a suitable matching with constant probability. Furthermore, since its running time is $O(n\log n)$ in expectation it runs in $O(n\log n)$ time with constant probability. Hence, through $O(\log n)$ calls to the algorithm and stopping early if the running time exceeds $O(n\log n)$ we can find a suitable matching $M_{S}$ with high probability in time $O(n\log^{2}n)$. The algorithm finally returns $M_{F}=M_{S}\cap E$

### v2

@@ 56    Delete-$M_{\texttt{adj}}$($(u,v)$)
Suppose that we maintain the matching $M_{\texttt{base}}$ as per Lemma˜6.6. Then with an additive overhead of $\tilde{O}\left(\delta n+\frac{B}{\delta}+\frac{n}{\delta B}\right)$ amortized update time, we can deterministically and explicitly maintain a maximal matching $M_{\texttt{adj}}\subseteq E_{\texttt{adj}}$ in $G_{\texttt{adj}}$.
We devote the rest of Section˜6.4 to the proof of Lemma˜6.14. We start by summarizing a few key properties of the adjunct graph $G_{\texttt{adj}}$ in the claim below.
@@
Consider any node $v\in V_{\texttt{adj}}\setminus V_{\texttt{dmg}}$. By part-(1) of Claim˜6.15, we have $v\notin V_{\texttt{sf}}$, and hence $v\notin V_{\texttt{hi}}$ (see Definition˜6.2). As the node-set $V$ is partitioned into the subsets $V_{\texttt{lo}},V_{\texttt{med}},V_{\texttt{hi}}$ (see Section˜6.2), it follows that $v\in V_{\texttt{lo}}\cup V_{\texttt{med}}$.
Since each phase lasts for $\delta n$ updates in $G$, we infer that $|E\setminus E_{\texttt{init}}|\leq\delta n$ (see Section˜6.2). Thus, the node $v$ can have at most $\delta n$ incident edges in $G_{\texttt{adj}}$ that are part of $E\setminus E_{\texttt{init}}$. Henceforth, we focus on bounding the number of incident edges on $v$ in $G_{res}$ that are part of $E\cap E_{\texttt{init}}$. We start by partitioning this concerned set of edges into three subsets, depending on their other endpoints:
$Z_{\texttt{lo}}$ $:=$ $\{(u,v)\in E_{\texttt{adj}}\cap E_{\texttt{init}}:u\in V_{\texttt{lo}}\cap V_{\texttt{adj}}\},$
$Z_{\texttt{med}}$ $:=$ $\{(u,v)\in E_{\texttt{adj}}\cap E_{\texttt{init}}:u\in V_{\texttt{med}}\cap V_{\texttt{adj}}\},$
$Z_{\texttt{hi}}$ $:=$ $\{(u,v)\in E_{\texttt{adj}}\cap E_{\texttt{init}}:u\in V_{\texttt{hi}}\cap V_{\texttt{adj}}\}.$
By part-(3) of Proposition˜6.1, we have $Z_{\texttt{lo}}\subseteq E(H_{\texttt{init}})$. Since $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it has maximum degree $B$ (see Definition˜2.1), and hence we get $|Z_{\texttt{lo}}|\leq B$. Next, part-(3) of Claim˜6.15 implies that $|Z_{\texttt{med}}|\leq 22\delta n=O(\delta n)$. Finally, since $V=V_{\texttt{sf}}\cup V_{\texttt{dmg}}$ (see Definition˜6.2), by part-(1) and part-(2), we get $|Z_{\texttt{hi}}|=O\left(\frac{n}{B}\right)$.
@@
At the start of a phase, we have $V_{\texttt{adj}}\cap V_{\texttt{hi}}=(V_{\texttt{adj}}\cap V_{\texttt{sf}})\cup(V_{\texttt{adj}}\cap V_{\texttt{dmg}})=\emptyset$, by Definition˜6.2 and part-(1) and part-(2) of Claim˜6.15. Accordingly, at this point in time, the node-set $V_{\texttt{adj}}$ is partitioned into two subsets: $V_{\texttt{adj}}\cap V_{\texttt{lo}}$ and $V_{\texttt{adj}}\cap V_{\texttt{med}}$.
By part-(3) of Claim˜6.15, there are at most $O(\delta n\cdot\delta n)$ edges in $E_{\texttt{adj}}$ whose both endpoints are in $V_{\texttt{adj}}\cap V_{\texttt{med}}$.
Next, consider any node $u\in V_{\texttt{adj}}\cap V_{\texttt{lo}}$. Since $V_{\texttt{adj}}\cap V_{\texttt{hi}}=\emptyset$ at this point in time, every edge in $E_{\texttt{adj}}$ that is incident on $u$ has its other endpoint in $V_{\texttt{lo}}\cup V_{\texttt{med}}$ (see Section˜6.2). Accordingly, by part-(3) of Proposition˜6.1, every edge in $E_{\texttt{adj}}\subseteq E_{\texttt{init}}$ that is incident on $u$ is part of $H_{\texttt{init}}$. Since $H_{\texttt{init}}$ is a $(B,(1-\epsilon)B)$-EDCS of $G_{\texttt{init}}$, it has maximum degree $B$ (see Definition˜2.1). Thus, the node $u\in V_{\texttt{adj}}\cup V_{\texttt{hi}}$ has degree at most $B$ in $G_{\texttt{adj}}$, at this point in time. We conclude that there are at most $O(nB)$ edges in $E_{\texttt{adj}}$ that have some endpoint in $V_{\texttt{adj}}\cap V_{\texttt{lo}}$.
@@
6.4.1 Proof of Lemma˜6.14.
At the start of a phase, we initialize $M_{\texttt{adj}}$ to be any arbitrary maximal matching in $G_{\texttt{adj}}$. By Corollary˜6.17, this takes $O(\delta n\cdot\delta n+Bn)$ time. Since the phase lasts for $\delta n$ updates in $G$, this step incurs an amortized update time of:
@@
Subsequently, to handle the updates during a phase, we use the following auxiliary data structures: Each node $v\in V_{\texttt{dmg}}$ maintains the set $\mathcal{F}_{\texttt{adj}}(v):=\left\{u\in V_{\texttt{adj}}\setminus V(M_{\texttt{adj}}):(u,v)\in E_{\texttt{adj}}\right\}$ of its free neighbors (in $G_{\texttt{adj}}$) under the matching $M_{\texttt{adj}}$. [^9] Maximality of $M_{\texttt{adj}}$ implies that $\mathcal{F}_{\texttt{adj}}(v)=\emptyset$ for all $v\in V_{\texttt{dmg}}\setminus V(M_{\texttt{adj}})$. Whenever a node $v$ moves from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$, we spend $\tilde{O}(n)$ time to initialize the set $\mathcal{F}_{\texttt{adj}}(v)$ as a balanced search tree. Within a phase, a node can move from $V_{\texttt{sf}}$ to $V_{\texttt{dmg}}$ at most once (see Definition˜6.2 and Corollary˜6.3). Thus, by part-(2) of Claim˜6.15, these initializations take $\tilde{O}\left(\frac{n}{B}\cdot n\right)=\tilde{O}\left(\frac{n^{2}}{B}\right)$ total time during a phase. As each phase lasts for $\delta n$ updates in $G$, this incurs an amortized update time of $\tilde{O}\left(\frac{n^{2}}{B\cdot\delta n}\right)=\tilde{O}\left(\frac{n}{\delta B}\right)$.
Since $V_{\texttt{adj}}:=V\setminus V(M_{\texttt{base}})$, Condition˜6.5 and Lemma˜6.6 imply that each edge update in $G$ leads to $O(1)$ node-updates in $G_{\texttt{adj}}$. Moreover, while maintaining the maximal matching $M_{\texttt{adj}}$ in $G_{\texttt{adj}}$, a node-update can be handled in $\tilde{O}\left(\delta n+B+\frac{n}{B}\right)$ worst-case time, for the following reason.
Suppose that we are trying to find a free neighbor (if it exists) of some node $v\in V_{\texttt{adj}}$. This can be done in $\mathrm{deg}_{G_{\texttt{adj}}}(v)=O\left(\delta n+B+\frac{n}{B}\right)$ time if $v\in V_{\texttt{adj}}\setminus V_{\texttt{dmg}}$ (see Corollary˜6.16), and in $\tilde{O}(1)$ time if $v\in V_{\texttt{dmg}}$ (using the set $\mathcal{F}_{\texttt{adj}}(v)$). Further, once we decide to match a node $u\in V_{\texttt{adj}}$ to one of its free neighbors, we can accordingly update all the relevant sets $\mathcal{F}_{\texttt{adj}}(v)$ in $\tilde{O}\left(\left|V_{\texttt{dmg}}\right|\right)=\tilde{O}\left(\frac{n}{B}\right)$ time (see part-(2) of Claim˜6.15).
@@
Lemma˜6.14 follows from (14) and (15).
@@
Theorem˜1.1 follows from Lemmas˜6.6 and 6.14.
To be a bit more specific, first recall that maintaining the EDCS $H$ incurs an update time of $O\left(\frac{n}{\epsilon B}\right)=O\left(\frac{n}{\delta B}\right)$, according to (7) and the first paragraph of Section˜6.1. Now, for the deterministic algorithm, part-(1) of Lemma˜6.6 incurs an amortized update time of $\tilde{O}\left(B\cdot n^{1/2}\cdot\delta^{-3/2}\right)$. Thus, by Lemma˜6.14, the overall amortized update time becomes
@@
Finally, for the randomized algorithm, part-(2) of Lemma˜6.6 incurs an amortized update time of $\tilde{O}(B\delta^{-1}+\delta^{-3})$. Thus, by Lemma˜6.14, the overall amortized update time becomes
@@ 7 Open Problems
The algorithm from Lemma˜5.1 is the main bottleneck in our dynamic maximal matching algorithm. In the static setting, we can deterministically find a perfect matching in an $m$-edge graph with $\gamma$-degree-gap with $\tilde{O}(m/\gamma)$ time, for example, using Dinic’s flow algorithm. Can we deterministically match this running time in the decremental setting? Such an algorithm would immediately imply a deterministic dynamic maximal matching algorithm with $\tilde{O}(n^{4/5})$ update time via our framework.
@@ Appendix A Convergence of Random Walks in Eulerian Expanders
In this section, we prove Theorem˜5.6 for completeness. We follow the notations from [RTV06]. We say that a matrix $M\in[0,1]^{V\times V}$ is a Markov chain with vertex set $V$ if $\sum_{v\in V}M(v,u)=1$ for every vertex $u$. Given a directed graph $G$, the random walk matrix $M_{G}$ of $G$ is the matrix where, for each entry $(v,u)$,
@@
For an Eulerian graph $G$, observe that the conductance of $M_{G}$ defined above coincides with the conductance of $G$ defined in Equation˜4. That is, we have
@@
With all lemmas above, we are now ready to prove Theorem˜5.6.
Proof of Theorem˜5.6.
@@
Let $\alpha$ be the distribution over $V$ where $\alpha(u)=1$ if $u=v$ and $\alpha(u)=0$ for $u\neq v$. Observe that $M_{G}^{k}\alpha$ is precisely the distribution over $V$ after making $k$ random walk steps in $G$ starting from vertex $v$. So $p$ is precisely the $t$-th entry of $M_{G}^{k}\alpha$. By Lemma˜A.3, we have that
@@
To bound $\lambda_{\pi}(M_{G})$, observe that $M_{G}(u,u)\geq 1/2$ for every $u$ because the number of self loops at $u$ is at least $\mathrm{deg}(u)/2$. So, by Lemma˜A.4 and Equation˜16, we have
@@
Appendix B Proof of Lemma˜3.1
@@
Now, we are ready to prove Lemma˜3.1.
@@
Note that when the algorithm of Theorem B.1 is called on $G^{\prime}$ in $O(|V^{\prime}|\log|V^{\prime}|)=O(n\log n)$ expected time it returns a matching $M$ such that every vertex in $V_{\kappa}$ is not matched by $M^{\prime}$ with probability at most $2/\Delta$. Therefore, the expected number of unmatched vertices in $V_{\kappa}$ under $M$ is at most $2|V_{\kappa}|/\Delta\leq 2n/\Delta\leq n\kappa/2$. Define a matching $M$ returned by the algorithm to be suitable if it doesn’t match at most $n\kappa$ vertices of $V_{\kappa}$. By Markov’s inequality we can argue that the algorithm of Theorem B.1 returns a suitable matching with constant probability. Furthermore, since its running time is $O(n\log n)$ in expectation it runs in $O(n\log n)$ time with constant probability. Hence, through $O(\log n)$ calls to the algorithm and stopping early if the running time exceeds $O(n\log n)$ we can find a suitable matching $M_{S}$ with high probability in time $O(n\log^{2}n)$. The algorithm finally returns $M_{F}=M_{S}\cap E$